OpenStudy (anonymous):

r(squared)-4r+c find the value of c that completes the square. how do you do this? :/

OpenStudy (angela210793):

hmmm....w8...

OpenStudy (angela210793):

c=2 (r-2)^2=r^2-2*r*2+2^2)=r^2-4r+4 :) i guess :):):)

OpenStudy (dumbcow):

(r+a)^2 = r^2 + 2ar +a^2 for this problem 2a = -4 a^2 = c 2a = -4 a = -2 c = (-2)^2 = 4

OpenStudy (angela210793):

oops yea c=2^2=4 :) sorry :)

OpenStudy (anonymous):

wait what does ^ stand for???

OpenStudy (angela210793):

@dumbcow u explain that please idk the word in english(english it's not my birth language) tht's y :)

OpenStudy (angela210793):

4 example 2^2=4 3^2=9 like squared or not...depends

OpenStudy (dumbcow):

its for exponents

OpenStudy (angela210793):

exactly!!! :)i just learnt a new word :) ^_^

OpenStudy (anonymous):

ohh got it got it, thanks guyys [: are you both students? high school ?

OpenStudy (angela210793):

yea 12th grade :S :S

OpenStudy (dumbcow):

im a student but way past high school, haha

OpenStudy (anonymous):

Haha I see. I am an upcoming junior .

OpenStudy (angela210793):

:)

OpenStudy (dumbcow):

nice, yeah this is a pretty good place to get help with hw and math questions lot of good math people here

OpenStudy (anonymous):

Yeah I can see that, I really like this website .